OpenStudy (anonymous):

if \[x = \frac{1}{3}\ \ \ and\ \ \ y = 7\] then find\[(1-x(\frac{y-1}{2}))^{x^3}\]

OpenStudy (jamesj):

why?

OpenStudy (anonymous):

My calculator gives me a different answers compared to what I get by hand.

OpenStudy (jamesj):

It's zero by inspection because y-1 = 6 => (y-1)/2 = 3 => x(y-1)/2 = 1

OpenStudy (anonymous):

wolfram gives a correct answer of 0

OpenStudy (anonymous):

\[\left(1-\frac{1}{3}\left(\frac{7-1}{2}\right)\right)^{(\frac{1}{3})^{3}}\]\[\left(1-\frac{1}{3}\left(\frac{6}{2}\right)\right)^{(\frac{1}{27})}\]\[\left(1-1\right)^{(\frac{1}{27})} \]\[0^{(\frac{1}{27})} = 0\]